A man does 4,650 J of work in the process of pushi
Answer #1

W = 4650 J
m = 3.4 * 10^3 Kg
Initial Velocity, vi = 0
Final Velocity, vf = ?
S = 21 m

(a)
Now,
Work done = Change in Kinetic Energy
4650 = 1/2*m*vf^2 - 1/2*m*vi^2
4650 = 1/2* 3.4 * 10^3 *vf^2- 0
vf = 1.65 m/s
Speed, v = 1.65 m/s

(b)
Work done = Force * distance
4650 = F* 21
F = 221.4 N
Horizontal Force, F = 221.4 N